# Document Valuation in LLM Summaries: A Cluster Shapley Approach

### v2

@@ 6 Implementation Details
Recall that the two key inputs to our document valuation framework are a summarization method $A(\cdot)$ and an evaluation method $v(\cdot)$. In $\S$6.1, we first introduce the RAG architecture, which describes how pre-trained LLMs can be augmented for search tasks with domain-relevant documents/articles. Next, in $\S$6.2 and $\S$6.3, we detail our summarization $A(\cdot)$ and evaluation $v(\cdot)$, respectively. In $\S$6.4, we discuss the challenges associated with stochasticity in summarization and evaluation in real settings and how we handle them. Finally, we discuss the implementation details for the exact Shapley and Cluster Shapley approaches in $\S$6.5 and $\S$$6.$
6.1 Introduction to RAG
AI search engines are designed to provide real-time, contextually relevant responses to user queries. A key technique behind many of these systems is RAG, which integrates pre-trained LLMs with information retrieval to enhance response accuracy and relevance (lewis2020retrieval). RAG addresses the limitations of static, pre-trained LLMs by incorporating new information from up-to-date, domain-relevant documents (which can be potentially proprietary to the firm). By grounding responses in reliable documents, RAG improves the relevance of AI-generated answers, reduces hallucinations, and mitigates the issue of outdated information that plagues static models (gao2023retrieval).
  • $\bullet$
  • $\bullet$
  • $\bullet$
To illustrate the impact of RAG integration on LLMs, we provide an example with a query: “Who won the Super Bowl this year?” in Figure 2. In this case, the ChatGPT application (powered by GPT-4o) leverages web browsing to retrieve up-to-date information that is not present in the model’s training data (GPT-4o has a knowledge cutoff of September 30, 2023). This example highlights how RAG enables LLMs to access and incorporate real-time information beyond their static training corpus. Note that the RAG system explicitly links to information sources in this example. This is standard in most RAG-based search systems since it allows users to verify response provenance and understand the basis of AI-generated content. Given the effectiveness and scalability of RAG architecture, it now forms the backbone of most of the modern AI-based search and summarization systems, e.g., Amazon’s Rufus (amazon_rufus), Google AI Search (google2024overview), and OpenAI’s ChatGPT Search (searchgpt).
@@ 6.1 Introduction to RAG
6.2 Summarization of Relevant Amazon Reviews via RAG
We now construct a RAG-based search to find relevant documents $S_{q}$ from Amazon reviews and summarization tool $A(q,S_{q})$ for any given query $q$. Figure 3 shows the overview of the four-step procedure.
@@ 6.2 Summarization of Relevant Amazon Reviews via RAG
    The pre-processing step consists of generating text embeddings for all $D$ reviews/documents associated with a product generated using OpenAI’s text-embedding-3-large model, which produces embeddings with a default size of 3072 dimensions. These embeddings are based on all the review text, including the title and the main content. In our analysis, we exclude reviews with fewer than 10 words, as they tend to be incomplete or uninformative. Our RAG architecture is agnostic to the exact source of embeddings, and it is possible to use alternative embedding models from open-source LLMs such as Llama, BERT, etc. However, recent research has shown that OpenAI embeddings tend to outperform the embeddings of such earlier models in discriminative tasks (ye2025lola); hence we use the OpenAI embeddings for our application.
@@
    We first process the user query to extract the key semantic information in it using a LLM (in our case GPT-4o-2024-08-06). The goal of this extraction is to identify the core meaning/consumer need expressed in the user’s query. For example, in Figure 3, the user’s query is, “I would like to know more details about the quality of the wireless controller.” Here, the key semantic information is, “quality of the wireless controller,” which is extracted for further processing.
@@
    Next, we use OpenAI’s text-embedding-3-large model to generate the embedding for the processed query. We denote the embedding of the query as $e_{q}$. For each review $i$ in the set of reviews $D$, we represent its embedding as $e_{i}$. We then calculate the cosine similarity between the query embedding $e_{q}$ and the review embedding $e_{i}$ for each review. The cosine similarity between the document embedding $e_{i}$ and the query embedding $e_{q}$ is defined as $\frac{e_{i}\cdot e_{q}}{\|e_{i}\|\|e_{q}\|}$, where $e_{i}\cdot e_{q}$ is the dot product of the embeddings, and $\|e_{i}\|$ and $\|e_{q}\|$ are their respective Euclidean norms. For a detailed explanation of cosine similarity and its application in text similarity tasks, see Chapter 6 of (schutze2008introduction). The cosine similarity of a pair of embedding vectors captures the extent to which the two vectors are similar, with higher values indicating greater similarity. Thus, a higher cosine similarity indicates that a given review $i$ has greater relevance to the query $q$. Next, we rank the cosine similarity scores of the query $q$ for all the $D$ reviews and retain the most relevant reviews. We choose $|S_{q}|=8$ as a fixed value for all queries, though in practice, the retrieval process can be more complex. [^8] For example, we can either retrieve the top few documents or apply a cutoff based on the cosine similarity of the embedding values—excluding those below a fixed threshold—or combine both approaches by first selecting the top few relevant documents and then filtering those that meet the similarity threshold.
@@
    We restrict ourselves to the eight most relevant reviews ($|S_{q}|=8$) from $D$ reviews/documents for three reasons. First, from a computational and monetary standpoint, including a large number of documents can be costly, as the LLM must process all associated tokens as context during generation. Second, including irrelevant or low-relevance documents may degrade summary quality, a concern supported by both anecdotal observations and prior literature on information overload (jacoby1974brand; eppler2004concept). Third, and most importantly, our choice reflects current industry practice. For example, (perplexity_citation_count) documented that Perplexity AI generated responses citing an average of 5.28 documents per query, and (google_citation_count) mentioned that “On average, (Google) AI overviews include 4-5 citations (the maximum is 33)” from analyzing 1 million overviews. In this context, summarizing the top 8 relevant sources aligns well with how real-world LLM summary applications present content today. That said, our proposed Cluster Shapley algorithm is not limited to small-scale applications. It can be readily scaled to larger document sets when needed. In $\S$7.3, we present additional results for settings where the number of relevant documents is large.
@@
    Given the user query ($q$) and the contextual information from the relevant reviews ($S_{q}$), the LLM (GPT-4o) produces a grounded summary by retrieving and synthesizing evidence from the provided documents. Although the LLM retains knowledge from its pretraining corpus, the response in this step is guided primarily by the contextual input delivered through the augmentation prompt (see Figure A6). The output summary cites supporting reviews in square brackets (e.g., [2]), enabling users to trace each statement back to specific source documents. If a review is not relevant, it is explicitly marked as such by the model, e.g., (“[4] is not related to the query”).
@@ 6.4 Stochasticity in Summarization and Evaluation
6.5 Exact Shapley Details and Costs
We now discuss the implementation details for both Exact Shapley and discuss the scalability and costs of wih this algorithm.
6.5.1 Exact Shapley Implementation and Example
For each product, we calculate the exact Shapley values using the formula in Equation (4). Table 2 presents the Shapley values for the top eight most relevant reviews in response to the query, “How is the quality of the wireless controller?” for the first product. Other reviews not contributing to this query receive a Shapley value of zero. In this example, we see that Review #3 has the highest Shapley value ($1.83$), as it directly compares the controller’s quality to other versions and emphasizes functionality, aligning well with the prompt’s emphasis on “Information Coverage” for quality details. Similarly, Review #7 ($1.61$) and Review #2 ($1.58$) score highly for addressing quality explicitly—#7 in a positive tone and #2 by highlighting durability compared to off-brand controllers. Review #5 ($1.44$) also performs well by underscoring the superior quality of the original controller versus knockoff brands. Review #4 ($1.25$) is somewhere in the middle, highlighting the good quality but without additional information relevant to the query. The lower-scoring reviews, including Review #1 ($0.59$) and Review #6 ($0.53$), just generally mention the great quality and cheap price, lacking specific details. Reviews #8 has the lowest value of $0.17$, as it emphasizes aspects like shipping and being good for gifts, which are less relevant to the query’s focus on controller quality, though the title mentions the “Quality”, which makes it relevant to the query.
@@ 6.5.1 Exact Shapley Implementation and Example
6.5.2 Computational Costs and Scalability
Next, we discuss the cost of implementing the exact Shapley algorithm. For each query with 8 relevant reviews, we must process $2^{8}-1=255$ distinct subsets, with each subset requiring a summarization and four evaluations to ensure reliable scoring. Our experiments indicate that processing a single query takes 15 minutes (on average) [^11] This time includes the full process for both summarization and evaluation of Python-based API calls, network latency, time to first token, and all computational overheads. Processing time depends significantly on the OpenAI API tier level; our experiments use Tier 2 access. and costs about \$1.30 in OpenAI API fees per query. As we can see, this can become prohibitive in both time and money as the number and variety of queries scale up. [^12] Batch processing, i.e., simultaneous API calls to OpenAI, can effectively reduce the processing time from 15 minutes to around 3.5 seconds by parallelizing the 255 summarizations and evaluations. However, the total computation time and cost remain unchanged. Alternatively, open-source LLMs for summarization and evaluation can further reduce both time and costs. For simplicity, we report the total computation time based on GPT-4o throughout the paper.
To illustrate the scalability challenge, consider Perplexity AI, a leading LLM-based search engine, which serves over 400 million queries per month (srinivas2024). Applying the exact Shapley algorithm to each query for this volume would imply over \$520 million in monthly compute costs (\$1.3 per query × 400 million queries). Switching to a lower-priced open-sourced LLM model such as Llama-3-8B-Instruct-Lite can reduce API costs by approximately 98.4% [^13] GPT-4o costs \$1.313 per million tokens compared to just \$0.021 per million tokens for Llama 3 (aimlapi2025). However, even with Llama 3, the cost to calculate exact Shapley values would still exceed \$8.3 million (\$520 million per month x (1- 98.4%)) per month, which is still too high compared to Perplexity’s annualized revenue of \$100 million as of March 2025 (srinivas2025). The scale of these computational expenses creates a significant gap between theoretical document valuation frameworks and practical implementation, highlighting the need for an efficient approximation algorithm like our proposed Cluster Shapley algorithm.
@@ 6.6 Cluster Shapley Implementation
After specifying the clustering diameter $\epsilon$, we apply Algorithm 2 to perform document clustering. Note that the distance matrix used in the algorithm is computed based on Equation (7). To obtain document embeddings, we extract query-relevant information from each comment individually, using the same LLM prompt as in the summarization step but applied to one document at a time. The computational cost of our adaptive clustering algorithm is negligible in our setting, especially when compared to the cost of LLM-based summarization and evaluation. We analyze and report the computation time of the clustering step using Algorithm 2 in Web Appendix §G.
Figure 4 presents the clustering results using Algorithm 2 for a sample query. In this instance, $\epsilon=0.05$, which yields six clusters. Increasing $\epsilon$ results in fewer clusters, further reducing computational cost, but may introduce higher approximation error. Even with six clusters, the computational complexity is significantly reduced—shrinking from $2^{8}-1=255$ to $2^{6}-1=63$, representing a fourfold improvement in efficiency.
@@
7 Results
We now present the results from the application of our Cluster Shapley algorithm to the Amazon review setting. In $\S$7.1, we first discuss a set of alternative Shapley approximation algorithms that can serve as benchmarks. Then, in $\S$7.2, we present the numerical results from our approach and present comparisons to the other benchmark algorithms. In $\S$7.3, we present a series of robustness checks and an extension.
7.1 Benchmark Algorithms
We now briefly summarize three widely used Shapley value approximation algorithms that serve as benchmarks against which we compare the performance of our proposed algorithm.
@@ 7.1 Benchmark Algorithms
    Monte Carlo: The Monte Carlo algorithm (or permutation sampling) is a popular approach for approximating Shapley values (mann1960values). This method randomly samples permutations from the $|S_{q}|!$ possible combinations of documents and then for each document $i$ and one permutation $P^{\pi}_{i}$, calculates its marginal contribution, i.e., $v(q,A(q,P^{\pi}_{i}\cup\{i\}))-v(q,A(q,P^{\pi}_{i}))$. Shapley value can then be approximated using the sample average of marginal contributions over all sampled permutations. As the number of permutation samples increases, the approximation error decreases, but the computational cost grows linearly. In our numerical experiments, we progressively increase the number of permutations to show the trade-off between accuracy and efficiency.
@@
    We briefly summarize the algorithm here and refer readers to ghorbani2019data for details. This algorithm randomly samples a permutation of reviews and sequentially calculates performance scores, $v$, by adding reviews in the permutation order. Since these scores are increasing, the algorithm truncates the computation by assigning zero marginal contributions to the remaining reviews when the gap between the current score and the maximum score (10 in our setting) is smaller than a pre-specified threshold, called performance tolerance. It basically means that when adding the remaining reviews, their marginal contributions are always smaller than this threshold. Thus, this algorithm simply assigns zero marginal contribution instead of calculating the negligible marginal value.
    The performance tolerance parameter, which governs the allowable change in Shapley values before truncation occurs, is tuned over a range of values – {0.1, 0.2, 0.3, 0.5, 0.7, 1, 2, 3} – to balance estimation accuracy and computational efficiency. Smaller values reduce the effectiveness of truncation—causing Truncated Monte Carlo to behave similarly to standard Monte Carlo—while larger values result in early truncation, degrading estimation accuracy. Based on the tuning, we select 0.5 for our main experiments. After fixing the performance tolerance at 0.5, we vary the number of sampled permutations to construct the efficient frontier, which illustrates the trade-off between computational cost and approximation error.
@@
    Kernel SHAP: Kernel SHAP is a model-agnostic approach to approximating Shapley values based on weighted least squares regression (lundberg2017unified). Our implementation uses Python’s SHAP package, which we adapt specifically for our LLM-based summarization task by implementing a custom mapping function between subset compositions and their corresponding summarization scores. The method employs KernelExplainer with an identity link function and L1 regularization to enhance numerical stability. We tested increasing numbers of samples to evaluate the performance of the method under different computational budgets. Kernel SHAP has been widely used across domains, including NLP for transformer-based model interpretation (kokalj2021bert), finance for credit risk analysis (fama2024explainable), healthcare for clinical decision support (li2022unified), and marketing for optimizing content engagement (kong2023neural). Recent work has pointed out some limitations of Kernel SHAP, especially when interactions are strong (gosiewska2019not; ragodos2024model). [^14] Note that the issues pointed out by these papers relate to the approximation process used in the Kernel SHAP algorithm, and not the Shapley concept itself. Nevertheless, given that it is one of the most widely adopted baseline algorithms used in practice, we include it in our evaluation for benchmarking purposes.
7.2 Numerical Comparison Results
Our numerical experiments include 48 test queries, designed as described in $\S$5. Each query comprises the eight most relevant reviews selected from the Amazon review dataset, forming the foundation for our comparative analysis of various Shapley value approximation algorithms.
To establish a stable evaluation baseline and reduce variance introduced by the summarization and evaluation steps, we standardize the process as follows: for each query, we generate a single summary for each subset (for all 255 possible subsets) and fix the evaluation score for each summary by averaging four evaluations, as detailed in $\S$6.3 and $\S$6.4. By fixing sample paths, we mitigate the inherent randomness in LLM outputs, ensuring consistent baseline measurements across different approximation methods.
@@ 7.2 Numerical Comparison Results
We now discuss the main takeaways from Figure 5. First, we see that Cluster Shapley achieves the best overall performance across all algorithms. Its error curve consistently lies below those of other algorithms, indicating that it achieves lower approximation error for a given computation cost or, equivalently, requires less computation to reach the same level of accuracy. Truncated Monte Carlo ranks second, benefiting from its early stopping mechanism, which limits unnecessary evaluations when performance plateaus. Second, Cluster Shapley’s advantage is particularly pronounced when the number of subsets is relatively small. In this regime, it achieves an MAE of around 0.2, whereas other methods exceed 0.4. This efficiency arises from clustering’s ability to capture core similarity patterns with relatively few clusters, which effectively reduces the dimensionality of the approximation problem. However, as the number of subsets increases (i.e., with smaller $\epsilon$), the advantage gradually diminishes. When the subset count exceeds 150, Cluster Shapley’s performance converges with other methods as single-review clusters emerge and the benefit of clustering is lost. Third, all algorithms show decreasing MAE as the number of subsets increases, consistent with theoretical expectations. However, their stability varies substantially. Monte Carlo and Truncated Monte Carlo exhibit noticeable fluctuations and wide confidence intervals due to their inherent sampling variability. In contrast, Cluster Shapley produces deterministic results across the computation range because we calculate the exact Shapley of clusters in Step 2. Kernel SHAP also uses sampling but reduces variance through its linear regression-based estimation, resulting in moderate stability.
Clustering Diameter ($\epsilon$) MAE MSE MAPE Cost Reduction
0.01 0.0381 0.0148 7.47% 23.01%
0.10 0.0507 0.0184 8.62% 26.67%
0.20 0.0913 0.0441 11.85% 40.00%
0.30 0.1617 0.1723 17.16% 62.39%
0.40 0.1972 0.1499 21.35% 73.61%
0.50 0.1908 0.1074 21.05% 81.52%
0.60 0.2152 0.1038 24.43% 91.62%
0.70 0.2305 0.1636 26.49% 98.63%
0.80 0.2259 0.2123 26.33% 99.13%
Table 3: Approximation error (averaged over all documents) and computation time reduction of Cluster Shapley under varying $\epsilon$. The last column, cost reduction, is calculated as the percentage reduction in the number of unique subsets used compared to all 255 subsets used by the exact Shapley. Note that when calculating MAPE, we add a small constant (0.1, approximately 10% of the mean Shapley value) to the denominator to prevent near-zero Shapley values from inflating the error.
We now present some additional results on the computational efficiency trade-offs of Cluster Shapley. Recall that the exact Shapley calculation requires evaluating all 255 possible unique subsets, with each subset evaluation taking an average of 3.5 seconds, resulting in a total computation time of approximately 15 minutes per query. Table  3 demonstrates how much this cost reduces when using Cluster Shapley for varying clustering diameters. We see that increasing the clustering diameter $\epsilon$ leads to greater computational savings but lower accuracy. Notably, at $\epsilon=0.20$, the algorithm achieves a 40% reduction in computation time while maintaining reasonable accuracy with an MAE of 0.0913 and MAPE of 11.85%. This represents an attractive balance point between efficiency and accuracy.

### v3

@@ 6 Implementation Details
The proposed document valuation framework is built on two key components: a summarization method $A(\cdot)$ and an evaluation method $v(\cdot)$, which are detailed in $\S$6.1 and $\S$6.2 respectively. Then, we discuss the implementation details for the exact Shapley and Cluster Shapley approaches in $\S$6.3 and $\S$6.4.
6.1 Summarization of Relevant Amazon Reviews via RAG
AI search engines are designed to provide real-time, contextually relevant responses to user queries. A key enabling technique behind many of these systems is retrieval-augmented generation (RAG), which integrates pre-trained LLMs with information retrieval modules to augment model outputs with domain-relevant documents (36). A detailed introduction to the RAG architecture is provided in Web Appendix §C.1.
To illustrate the impact of RAG integration on LLMs, we provide an example with a query: “Who won the Super Bowl this year?” in Figure 2. In this case, the ChatGPT application (powered by GPT-4o) leverages web browsing to retrieve up-to-date information that is not present in the model’s training data (GPT-4o has a knowledge cutoff of September 30, 2023). This example highlights how RAG enables LLMs to access and incorporate real-time information beyond their static training corpus. Note that the RAG system explicitly links to information sources in this example. This is standard in most RAG-based search systems since it allows users to verify response provenance and understand the basis of AI-generated content. Given the effectiveness and scalability of RAG architecture, it now forms the backbone of most of the modern AI-based search and summarization systems, e.g., Amazon’s Rufus (43), Google AI Search (55), and OpenAI’s ChatGPT Search (47).
@@ 6.1 Summarization of Relevant Amazon Reviews via RAG
Next, we construct a RAG-based search to find relevant documents $S_{q}$ from Amazon reviews and summarization tool $A(q,S_{q})$ for any given query $q$. Figure 3 shows the overview of the four-step procedure.
@@
    The pre-processing step consists of generating text embeddings for all $D$ reviews/documents associated with a product generated using OpenAI’s text-embedding-3-large model, which produces embeddings with a default size of 3072 dimensions. These embeddings are based on all the review text, including the title and the main content. In our analysis, we exclude reviews with fewer than 10 words, as they tend to be incomplete or uninformative. Our RAG architecture is agnostic to the exact source of embeddings, and it is possible to use alternative embedding models from open-source LLMs such as Llama, BERT, etc. However, recent research has shown that OpenAI embeddings tend to outperform the embeddings of such earlier models in discriminative tasks (69); hence we use the OpenAI embeddings for our application.
@@
    We first process the user query to extract the key semantic information in it using an LLM (in our case GPT-4o-2024-08-06). The goal of this extraction is to identify the core meaning/consumer need expressed in the user’s query. For example, in Figure 3, the user’s query is, “I would like to know more details about the quality of the wireless controller.” Here, the key semantic information is, “quality of the wireless controller,” which is extracted for further processing.
@@
    Next, we use OpenAI’s text-embedding-3-large model to generate the embedding for the processed query. We denote the embedding of the query as $e_{q}$. For each review $i$ in the set of reviews $D$, we represent its embedding as $e_{i}$. We then calculate the cosine similarity between the query embedding $e_{q}$ and the review embedding $e_{i}$ for each review. The cosine similarity between the document embedding $e_{i}$ and the query embedding $e_{q}$ is defined as $\frac{e_{i}\cdot e_{q}}{\|e_{i}\|\|e_{q}\|}$, where $e_{i}\cdot e_{q}$ is the dot product of the embeddings, and $\|e_{i}\|$ and $\|e_{q}\|$ are their respective Euclidean norms. For a detailed explanation of cosine similarity and its application in text similarity tasks, see Chapter 6 of (60). The cosine similarity of a pair of embedding vectors captures the extent to which the two vectors are similar, with higher values indicating greater similarity. Thus, a higher cosine similarity indicates that a given review $i$ has greater relevance to the query $q$. Next, we rank the cosine similarity scores of the query $q$ for all the $D$ reviews and retain the most relevant reviews.
@@
    We restrict ourselves to the eight most relevant reviews ($|S_{q}|=8$) from $D$ for all queries, though in practice, the retrieval process can be more complex. [^9] We can either retrieve the top few documents or apply a cutoff based on the cosine similarity of the embedding values-excluding those below a fixed threshold-or combine both approaches by first selecting the top few relevant documents and then filtering those that meet the similarity threshold. Our choice reflects current industry practice. For example, (13) documented that Perplexity AI generated responses citing an average of 5.28 documents per query, and (26) mentioned that “On average, (Google) AI overviews include 4-5 citations (the maximum is 33)” from analyzing 1 million overviews. In this context, summarizing the top 8 relevant sources aligns well with how real-world LLM summary applications present content today. That said, our proposed Cluster Shapley algorithm is not limited to small-scale applications. It can be readily scaled to larger document sets when needed.
@@
[…]
@@ 6.2 Evaluation of Summarized Amazon Reviews
6.3 Exact Shapley Value Calculation
We now discuss the implementation details for both the exact Shapley value calculation and discuss the scalability and costs of this algorithm.
For each product, we calculate the exact Shapley values using the formula in Equation (4). Table 2 presents the Shapley values for the top eight most relevant reviews in response to the query, “How is the quality of the wireless controller?” for the first product. Other reviews not contributing to this query receive a Shapley value of zero. In this example, we see that Review #3 has the highest Shapley value ($1.83$), as it directly compares the controller’s quality to other versions and emphasizes functionality, aligning well with the prompt’s emphasis on “Information Coverage” for quality details. Similarly, Review #7 ($1.61$) and Review #2 ($1.58$) score highly for addressing quality explicitly-#7 in a positive tone and #2 by highlighting durability compared to off-brand controllers. Review #5 ($1.44$) also performs well by underscoring the superior quality of the original controller versus knockoff brands. Review #4 ($1.25$) is somewhere in the middle, highlighting the good quality but without additional information relevant to the query. The lower-scoring reviews, including Review #1 ($0.59$) and Review #6 ($0.53$), just generally mention the great quality and cheap price, lacking specific details. Reviews #8 has the lowest value of $0.17$, as it emphasizes aspects like shipping and being good for gifts, which are less relevant to the query’s focus on controller quality, though the title mentions the “Quality”, which makes it relevant to the query.
@@ 6.3 Exact Shapley Value Calculation
Next, we discuss the cost of implementing the exact Shapley calculation. For each query with 8 relevant reviews, we must process $2^{8}-1=255$ distinct subsets, with each subset requiring a summarization and four evaluations to ensure reliable scoring. We use OpenAI’s API to process both summarization and evaluation with GPT-4o, and it costs about \$1.30 in OpenAI API fees per query.
To illustrate the scalability challenge, consider Perplexity AI, a leading LLM-based search engine, which serves over 400 million queries per month (64). Applying the exact Shapley algorithm to each query for this volume would imply over \$520 million in monthly compute costs (\$1.3 per query × 400 million queries). Switching to a lower-priced open-sourced LLM model such as Llama-3-8B-Instruct-Lite can reduce API costs by approximately 98.4% [^12] GPT-4o costs \$1.313 per million tokens compared to just \$0.021 per million tokens for Llama 3 (1). However, even with Llama 3, the cost to calculate exact Shapley values would still exceed \$8.3 million (\$520 million per month x (1- 98.4%)) per month, which is still too high compared to Perplexity’s annualized revenue of \$100 million as of March 2025 (65). The scale of these computational expenses creates a significant gap between theoretical document valuation frameworks and practical implementation, highlighting the need for an efficient approximation algorithm like our proposed Cluster Shapley algorithm.
@@ 6.4 Cluster Shapley Implementation
After specifying the clustering diameter $\epsilon$, we apply Algorithm 3 to perform document clustering. Note that the distance matrix used in the algorithm is computed based on Equation (7). To obtain document embeddings, we extract query-relevant information from each comment individually, using the same LLM prompt as in the summarization step but applied to one document at a time. The computational cost of our adaptive clustering algorithm is negligible in our setting, especially when compared to the cost of LLM-based summarization and evaluation. We analyze and report the computation time of the clustering step using Algorithm 3 in Web Appendix §C.5.
Figure 4 presents the clustering results using Algorithm 3 for a sample query. In this instance, $\epsilon=0.05$, which yields six clusters. Increasing $\epsilon$ results in fewer clusters, further reducing computational cost, but may introduce higher approximation error. Even with six clusters, the computational complexity is significantly reduced-shrinking from $2^{8}-1=255$ to $2^{6}-1=63$, representing a fourfold improvement in efficiency.
@@
7 Numerical Results
We now present the results from the application of our Cluster Shapley algorithm to the Amazon review setting. First, in $\S$7.1, we report the numerical results of our proposed Cluster Shapley approach and compare its performance with other three off-the-shelf Shapley approximation algorithms. Next in $\S$7.2, we further contrast Cluster Shapley with the simple attribution rules (with minimal computational cost) and discusses the corresponding findings. In $\S$7.3, we present a series of robustness checks and extensions.
7.1 Comparison with Off-the-shelf Shapley Approximation Algorithms
We now introduce three widely used Shapley value approximation algorithms that serve as off-the-shelf benchmarks and compare the performance of our proposed algorithm against them.
@@ 7.1 Comparison with Off-the-shelf Shapley Approximation Algorithms
    Monte Carlo: The Monte Carlo algorithm (or permutation sampling) is a popular approach for approximating Shapley values (42). This method randomly samples permutations from the $|S_{q}|!$ possible combinations of documents and then for each document $i$ and one permutation $P^{\pi}_{i}$, calculates its marginal contribution, i.e., $v(q,A(q,P^{\pi}_{i}\cup\{i\}))-v(q,A(q,P^{\pi}_{i}))$. Shapley value can then be approximated using the sample average of marginal contributions over all sampled permutations. As the number of permutation samples increases, the approximation error decreases, but the computational cost grows linearly. In our numerical experiments, we progressively increase the number of permutations to show the trade-off between accuracy and efficiency.
@@
    We briefly summarize the algorithm here and refer readers to Ghorbani and Zou (20) for details. This algorithm randomly samples a permutation of reviews and sequentially calculates performance scores, $v$, by adding reviews in the permutation order. Since these scores are increasing, the algorithm truncates the computation by assigning zero marginal contributions to the remaining reviews when the gap between the current score and the maximum score (10 in our setting) is smaller than a pre-specified threshold, called performance tolerance. It basically means that when adding the remaining reviews, their marginal contributions are always smaller than this threshold. Thus, this algorithm simply assigns zero marginal contribution instead of calculating the negligible marginal value.
    The performance tolerance parameter, which governs the allowable change in Shapley values before truncation occurs, is tuned over a range of values – {0.1, 0.2, 0.3, 0.5, 0.7, 1, 2, 3} – to balance estimation accuracy and computational efficiency. Smaller values reduce the effectiveness of truncation-causing Truncated Monte Carlo to behave similarly to standard Monte Carlo-while larger values result in early truncation, degrading estimation accuracy. Based on the tuning, we select 0.5 for our main experiments. After fixing the performance tolerance at 0.5, we vary the number of sampled permutations to construct the efficient frontier, which illustrates the trade-off between computational cost and approximation error.
@@
    Kernel SHAP: Kernel SHAP is a model-agnostic approach to approximating Shapley values based on weighted least squares regression (40). Our implementation uses Python’s SHAP package, which we adapt specifically for our LLM-based summarization task by implementing a custom mapping function between subset compositions and their corresponding summarization scores. The method employs KernelExplainer with an identity link function and L1 regularization to enhance numerical stability. We tested increasing numbers of samples to evaluate the performance of the method under different computational budgets. Kernel SHAP has been widely used across domains, including NLP for transformer-based model interpretation (33), finance for credit risk analysis (16), healthcare for clinical decision support (37), and marketing for optimizing content engagement (34). Recent work has pointed out some limitations of Kernel SHAP, especially when interactions are strong (23; 53). [^13] Note that the issues pointed out by these papers relate to the approximation process used in the Kernel SHAP algorithm, and not the Shapley concept itself. Nevertheless, given that it is one of the most widely used Shapley approximations used in practice, we include it in our evaluation for benchmarking purposes.
We now present the results from the empirical exercise comparing the performance of our algorithm with the above three algorithms. Our experiments include the 48 test queries described in $\S$5. Each query comprises the eight most relevant reviews selected from the Amazon review dataset, forming the foundation for our comparative analysis of various Shapley value approximation algorithms.
To establish a stable evaluation baseline and reduce variance introduced by the summarization and evaluation steps, we standardize the process as follows: for each query, we generate a single summary for each subset (for all 255 possible subsets) and fix the evaluation score for each summary by averaging four evaluations, as detailed in $\S$6.2. By fixing sample paths, we mitigate the inherent randomness in LLM outputs, ensuring consistent baseline measurements across different approximation methods.
@@
We now discuss the main takeaways from Figure 5. First, we see that Cluster Shapley achieves the best overall performance across all algorithms. Its error curve consistently lies below those of other algorithms, indicating that it achieves lower approximation error for a given computation cost or, equivalently, requires less computation to reach the same level of accuracy. Truncated Monte Carlo ranks second, benefiting from its early stopping mechanism, which limits unnecessary evaluations when performance plateaus. Second, Cluster Shapley’s advantage is particularly pronounced when the computation budget is small, i.e., the number of unique subsets is relatively small. In this regime, it achieves an MAE of around 0.2, whereas other methods exceed 0.4. This efficiency arises from clustering’s ability to capture core similarity patterns with relatively few clusters, which effectively reduces the dimensionality of the approximation problem. However, as the number of subsets increases (i.e., with smaller $\epsilon$), the advantage gradually diminishes. When the subset count exceeds 150, Cluster Shapley’s performance converges with other methods as single-review clusters emerge and the benefit of clustering is lost. Third, all algorithms show decreasing MAE as the number of subsets increases, consistent with theoretical expectations. However, their stability varies substantially. Monte Carlo and Truncated Monte Carlo exhibit noticeable fluctuations and wide confidence intervals due to their inherent sampling variability. In contrast, Cluster Shapley produces deterministic results across the computation range because we calculate the exact Shapley of clusters in Step 2. Kernel SHAP also uses sampling but reduces variance through its linear regression-based estimation, resulting in moderate stability.